# Motives of nullcones of quiver representations

### v1

@@ 1. Introduction
Our first result gives an explicit formula for the motives, as a recursion formula (Corollary 3.1), a generating function identitfy (Corollary 3.2), and as an explicit sum formula (Corollary 3.3). These formulas are derived from a geometric construction essentially using the stratification of nilpotent quiver representations by their socle type, inspired by methods of [11] (Section 2).
@@
Our second result (sections 4, 5) interprets the Hesselink stratification of [8] as an identity of motives, first yielding a positive summation formula (Theorem 5.2) for the motive of the nullcone, and culminating in a product formula relating the generation function of motives of nullcones to generating functions of semistable representations of level quivers (Theorem 5.7), reminiscent of wall-crossing formulas such as [10, Theorem 2.4].
@@
Acknowledgments: The second named author would like to S. Mozgovoy for stimulating discussions. The work of both authors is supported by the DFG GRK 2240 “Algebro-Geometric Methods in Algebra, Arithmetic and Topology”.
@@ 2.1. Quiver representations and their nullcone
Let $N_{\mathbf{d}}$ be the nullcone for this action. By the Hilbert criterion, it can be characterized by the following equivalent conditions:
@@
The following are equivalent for a representation $V\in R_{\mathbf{d}}$:
@@ 2.2. Motives
$[{\rm GL}_{n}(\mathbb{C})]=\mathbb{L}^{n(n-1)/2}(\mathbb{L})_{n},\;[{\rm Gr}_{k}^{n}]=\left[{n\atop k}\right]_{\mathbb{L}}:=\frac{(\mathbb{L})_{n}}{(\mathbb{L})_{k}\cdot(\mathbb{L})_{n-k}}.$
@@ 2.3. The recursive formula
Assume ${\bf d}\not=0$. For a vertex $i\in Q_{0}$ and a representation $V\in N_{\bf d}$ given by a tuple $(V_{\alpha})_{\alpha})$, we consider the map
@@
We note the following identity of Gaussian binomial coefficients for all $m\geq 0$ (see for example [11, (3.14)]):
@@
Finally, substituting ${\bf f}={\bf d}-{\bf e}$, the claimed formula follows, finishing the proof.
@@ 3. Applications and examples
The recursion follows from Theorem 2.2 just by making the motive of $G_{\bf e}$ explicit. Defining polynomials $n_{\bf d}$ by the same recursion with $\mathbb{L}$ replaced by $x$, and starting with $n_{0}=1$, we obviously have $n_{\bf d}\in\mathbb{Z}[x^{\pm 1}]$, and $n_{\bf d}(\mathbb{L})=[N_{\bf d}]$. By the definition of the Euler form, the $\mathbb{L}$-exponent in the recursion is always nonnegative, thus we have $n_{\bf d}(x)\in\mathbb{Z}[x]$.
@@
which equals 00 for ${\bf d}\not=0$ by Theorem 2.2.
@@
=$\sum_{s\geq 0}\sum_{{\bf d}^{1},\ldots,{\bf d}^{s}\not=0}\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\frac{t^{\bf d}}{(\mathbb{L})_{{\bf d}^{1}}\cdot\ldots\cdot(\mathbb{L})_{{\bf d}^{s}}}.$
Comparing coefficients of $t^{\bf d}$ on both sides and rewriting $[G_{\bf d}]$ using the Pochhammer symbol, we arrive at the claimed formula.
@@
fulfills $x^{\bf d}\cdot x^{\bf e}=x^{{\bf d}+{\bf e}}$. We extend the base ring of motives by adjoinging a square root of $\mathbb{L}$. This allows us to define motivic Donaldson–Thomas invariants by the identity
@@
Interchanging $\mathbb{L}$ and $\mathbb{L}^{-1}$, we thus have
@@
Combining these identities, the claim follows.
@@
$\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i,j}\sum_{k}r_{ij}\delta_{k\in K_{i}}\delta_{l\in K_{j}})=$
$=\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{k}\sum_{i}r_{ii})=\sum_{i}r_{ii}{d_{i}\choose 2}+\sum_{i<j}r_{ij}d_{i}d_{j}.$
All of the corresponding summands have the same sign $(-1)^{\dim{\bf d}}$, thus there are no cancellations in the summation of these top degree terms. The number of these summands equals the number of decompositions of $\{1,\ldots,\dim{\bf d}\}$ into parts $K_{i}$ for $i\in Q_{0}$, which is the quotient of factorials of the claimed formula. Adding the contribution $\sum_{i}{d_{i}\choose 2}$ of the prefactor in the the above formula for $n_{\bf d}(x)$ to the degree, we arrive at the claimed formula for the leading term.
Remark: A similar proof appears in [12, Theorem 4.1]. Refining the argument there shows that, in case $Q$ is symmetric, the variety $N_{\bf d}$ is equidimensional of dimension
@@
    $t_{i_{n}}^{d_{i_{1}}}\cdot\ldots\cdot t_{i_{1}}^{d_{i_{n}}}=\mathbb{L}^{-\sum_{k}d_{i_{k}}(d_{i_{k}}-1)/2}t^{\bf d}.$
@@
    $\sum_{\bf d}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}}=\prod_{k=1}^{n}\sum_{d\geq 0}\frac{\mathbb{L}^{d(d-1)/2}t_{n+1-k}^{d}}{(\mathbb{L})_{d}}$
@@
    $(\sum_{\bf d}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}})^{-1}=\prod_{k=1}^{n}\sum_{d\geq 0}\frac{(-t_{k})^{d}}{(\mathbb{L})_{d}}=$
@@
    It is easy to prove that, for general $d$, a monomial $\mathbb{L}^{a}$ appears in $[N_{d}]$ with non-zero coefficient if and only if $a=km+l$ for $d-1\leq k\leq{d\choose 2}$ and (${d\choose 2}-k\leq l\leq{d\choose 2}$.
@@ 4.1. Introduction
$I_{v}=[\sum_{i=0}^{v-1}d_{i}+1\ldots\sum_{i=1}^{v}d_{i}]~{}.$
Note that we need to fix an enumeration of the vertices of $Q$, but the following constructions will be independent of this choice.
@@
If $R$ is unstable, there is a unique coweight $s(R)\subset\mathbb{Q}^{d}$ with this property and such that the norm $|s(R)|=s_{1}^{2}+\dots+s_{d}^{2}$ is minimal. We will call such a coweight minimal.
@@
If $R$ is a saturated subset of $\Pi$, we have a corresponding saturated subspace $X_{R}$ of $N_{\mathbf{d}}(Q)$ by taking
@@
By [5] there is a bijection between the $G_{\mathbf{d}}$-conjugacy classes of saturated subspaces of $N_{\mathbf{d}}(Q)$ and the conjugacy classes of saturated subsets $R\subset\Pi$ under the action of the Weyl group $S_{\mathbf{d}}=S_{d_{1}}\times\dots\times S_{d_{n}}$ of $G_{\mathbf{d}}$. This correspondence assigns to $R$ the subspace $X_{R}$, and to a saturated subspace the set of nonzero weights of its elements. Hence, the number of Hesselink strata of $N_{\mathbf{d}}(Q)$ is at most the number of conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group.
We construct a bijection between the conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group and coweights satisfying certain additional properties. Consider a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ for $G_{\mathbf{d}}$. Then there is a unique partition of $s$ into a minimal number of pairwise disjoint sets $J$, called segments, such that the elements of each segment form an ordered sequence increasing by 00 or $1$ in each step. This induces a partition of $[1\ldots d]$ whose sets will be denoted by $\tilde{J}$. After introducing this notation we can now state definitions for the required conditions on coweights.
@@
The following lemma characterizes the property of being minimal and substantially improves [8], where only one direction is proven; see also [5].
@@
A coweight $s=(s_{1},\dots,s_{n})\in\mathbb{Q}^{d}$ is balanced if and only if it is minimal.
@@
Suppose $\mu=(\mu_{1},\dots,\mu_{n})$ is not balanced. Then there exists a segment $\tilde{J}$ such that $\sum_{i\in\tilde{J}}\mu_{i}\neq 0$. By definition of the segments, for every $\mu_{j}\in J$ and $\mu_{k}\notin J$ we have $|\mu_{j}-\mu_{k}|<1$ or $|\mu_{j}-\mu_{k}|>1$. Consequently we can choose an $\epsilon>0$ such that these inequalities also hold for $\mu_{j}+\epsilon$ and $\mu_{j}-\epsilon$. By shifting all entries of $J$ by $\epsilon$ or $-\epsilon$, the distance between two entries in the segment does not change. In this way we construct another string $\mu^{\prime}$ such that $R(\mu)=R(\mu^{\prime})$. Without loss of generality let $\sum_{i\in\tilde{J}}\mu_{i}=\alpha>0$ and $\mu^{\prime}=\mu-\epsilon$. We obtain:
$|\mu^{\prime}|=\sum_{i=1}^{n}(\mu^{\prime}_{i})^{2}=\sum_{i=1}^{n}(\mu-\epsilon)^{2}=\sum_{i=1}^{n}\mu^{2}-2\mu\epsilon+\epsilon^{2}=|\mu|-2\alpha\epsilon+n\epsilon^{2}~{}.$
@@
By choosing $\epsilon<\frac{2\alpha}{n}$, we obtain $\mu^{\prime}<\mu$. Accordingly, $\mu$ is not minimal. The case $\sum_{i\in\tilde{J}}\mu_{i}<0$ follows analogously with $\mu^{\prime}_{i}=\mu+\epsilon$.
Suppose now $\mu=(\mu_{1},\dots,\mu_{n})$ is balanced, but not minimal. Therefore, there exists a $\mu^{\prime}$ with $|\mu|>|\mu^{\prime}|$ and $R(\mu)\subset R(\mu^{\prime})$. We consider an arbitrary segment $J$ of $\mu$ of the form
$s=(\alpha,\dots,\alpha,\alpha+1,\dots,\alpha+1,\dots,\alpha+k,\dots,\alpha+k)~{}.$
We denote by $a_{i}$ the multiplicity of $\alpha+i$ in $s$ and define $a=\sum_{i=1}^{k}a_{i}$.
We now consider the corresponding elements in $\mu^{\prime}$, that is, we single out the elements of $\mu^{\prime}$ in the same positions as the elements of $J$ in $\mu$, and denote by $t$ the tuple listing them in weakly ascending order. Since $R(\mu)\subset R(\mu^{\prime})$, the tuple $t$ must be of the form
$(\beta,\beta+\tilde{\epsilon}_{1},\beta+\tilde{\epsilon}_{2},\dots,\beta+\tilde{\epsilon}_{a_{0}},\beta+1+\tilde{\epsilon}_{a_{0}+1},\dots,\beta+1+\tilde{\epsilon}_{a_{0}+a_{1}},\dots,\beta+k+\tilde{\epsilon}_{a})$
with $\tilde{\epsilon}_{i}=\sum_{j=1}^{i}\epsilon_{j}$ and $\epsilon_{j}\geq 0$ for $1\leq j\leq a$.
Without loss of generality we can shift $t$ to $t^{\prime}$ by adding a constant to each entry, such that
$\sum_{i=1}^{a}t^{\prime}_{i}=0~{},$
since, similarly to the first part of the proof, we have $|t|\geq|t^{\prime}|$.
Now we consider the two cases $\alpha\leq\beta^{\prime}$ and $\alpha>\beta^{\prime}$.
If $\alpha\leq\beta^{\prime}$, then obviously $s_{i}\leq t^{\prime}_{i}$ for all $1\leq i\leq a$, where $s_{i}$ and $t^{\prime}_{i}$ denote the $i$-th element of $s$ and $t^{\prime}$, respectively. Since $\sum_{i=1}^{a}t^{\prime}_{i}=0$ and $\sum_{i=1}^{a}s_{i}=0$, it follows that $t^{\prime}_{i}=s_{i}$. Thus $|\mu^{\prime}|\geq|\mu|$ since we chose an arbitrary segment. This is a contradiction to the assumption $|\mu|>|\mu^{\prime}|$.
If $\alpha>\beta^{\prime}$, there exists an entry $t^{\prime}_{k}$ such that $t^{\prime}_{k}\geq s_{k}$. We write $I=[1,\dots,a]$ and define $I^{s}_{n}=[1,\dots,m]$ and $I^{s}_{p}=[m+1,\dots,a]$, where $m$ is given by $s_{m}\leq 0$ and $s_{m+1}>0$. Similarly, we define $I_{n}^{t^{\prime}}$ and $I_{p}^{t^{\prime}}$. Note that
(1) $\sum_{i\in I_{n}}(-s_{i})=\sum_{i\in I_{p}}s_{i}=\frac{1}{2}|s_{i}|~{}.$
A similar statement holds for $I^{t^{\prime}}_{n}$ and $I^{t^{\prime}}_{p}$.
Suppose $t^{\prime}_{k}\geq 0$. Then $s_{i}>t^{\prime}_{i}$ for all $i\in I_{n}^{s}$ and consequently $|s_{i}|<|t^{\prime}_{i}|$. This yields
$\sum_{i\in I^{s}_{n}}|s_{i}|^{2}<\sum_{i\in I^{s}_{n}}|t^{\prime}_{i}|^{2}\leq\sum_{i\in I^{t^{\prime}}_{n}}|t^{\prime}_{i}|^{2}~{}.$
With (1) it follows that $|s|\geq|t^{\prime}|$, contrary to the assumption that $|\mu|>|\mu^{\prime}|$, since we started with an arbitrary segment and $|t^{\prime}|\leq|t|$. The case $t^{\prime}_{k}<0$ follows similarly, proving that a balanced string is automatically minimal. ∎
@@
We can thus describe the index set of the Hesselink strata using balanced and dominant coweights. The problem of detecting which strata are non-empty remains. To this effect, for every coweight $s$, we define a tuple $(Q_{s},{\bf d}_{s},\theta_{s})$, consisting of a quiver, a dimension vector and a stability, as follows:
@@
$\{(v,j)\;|\;v\in Q_{0}\text{ and }1\leq j\leq k_{i}+1\}~{},$
where $p_{i}+k_{i}$ is the last entry in $J_{i}$. In $Q_{i}$ the number of arrows from $(v,k)$ to $(v^{\prime},k+1)$ equals the number of arrows from $v$ to $v^{\prime}$ in $Q$. The dimension vector ${\bf d}_{i}$ of $Q_{i}$ is given by $(a_{i,k-1}^{v})_{v,k}$.
@@
A stability on $Q_{s}$ is defined by $\theta_{s}=(m_{v,j}^{(i)})_{i,v,j}$, where $m_{v,j}^{(i)}=-z\cdot(p_{i}+j)$ and $z$ is the least common multiple of all numerators of rational numbers $p_{i}$ with $1\leq i\leq u$.
@@
The following theorem summarizes the main properties relating the representation spaces of level quivers and the Hesselink stratification:
@@
Consider $|Q_{1}|$-tuples of $d\times d$-matrices. Then $V_{s}$ is given as the affine space of such tuples, where the matrix corresponding to an arrow $\alpha:v\rightarrow v^{\prime}$ has arbitrary entries at a position $(j,i)$ if $i\in I_{v}$, $j\in I_{v^{\prime}}$ and $s_{j}-s_{i}\geq 1$, and zero entries everywhere else (compare with Definition 4.1). The group $P_{s}$ is given by a single $d\times d$-matrix with zero entries everywhere except possibly at a position $(j,i)$ if $i,j\in I_{v}$ for some vertex $v\in Q_{0}$ and $s_{j}-s_{i}\geq 0$. For the affine subspace $B_{s}\subset V_{s}$ obtained by additionally requiring zero entries at all positions $(j,i)$ except the ones where $s_{j}-s_{i}=1$, we have

### v2

@@ 1. Introduction
Our first result gives an explicit formula for the motives, as a recursion formula (Corollary 3.1), a generating function identity (Corollary 3.2), and as an explicit sum formula (Corollary 3.3). These formulas are derived from a geometric construction essentially using the stratification of nilpotent quiver representations by their socle type, inspired by methods of [12] (Section 2).
@@
Our second result (sections 4, 5) interprets the Hesselink stratification of [8] as an identity of motives, first yielding a positive summation formula (Theorem 5.2) for the motive of the nullcone, and culminating in a product formula relating the generating function of motives of nullcones to generating functions of semistable representations of level quivers (Theorem 5.7), reminiscent of wall-crossing formulas such as [11, Theorem 2.4].
@@
Acknowledgments: The second named author would like to thank S. Mozgovoy for stimulating discussions. The work of both authors is supported by the DFG GRK 2240 “Algebro-Geometric Methods in Algebra, Arithmetic and Topology”.
@@ 2.1. Quiver representations and their nullcone
Let $N_{\mathbf{d}}$ be the nullcone for this action. By the its definition and the Hilbert criterion, it can be characterized by the following equivalent conditions:
@@
[8, Section 1.4] The following are equivalent for a representation $V\in R_{\mathbf{d}}$:
@@ 2.2. Motives
$[{\rm GL}_{n}(\mathbb{C})]=(-1)^{n}\mathbb{L}^{n(n-1)/2}(\mathbb{L})_{n},\;[{\rm Gr}_{k}^{n}]=\left[{n\atop k}\right]_{\mathbb{L}}:=\frac{(\mathbb{L})_{n}}{(\mathbb{L})_{k}\cdot(\mathbb{L})_{n-k}}.$
@@ 2.3. The recursive formula
Assume ${\bf d}\not=0$. For a vertex $i\in Q_{0}$ and a representation $V\in N_{\bf d}$ given by a tuple $(V_{\alpha})_{\alpha}$, we consider the map
@@
We note the following identity of Gaussian binomial coefficients for all $m\geq 0$ (see for example [12, (3.14)]):
@@
Finally, substituting ${\bf f}={\bf d}-{\bf e}$, the claimed formula follows, finishing the proof.∎
@@ 3. Applications and examples
The recursion follows from Theorem 2.2 just by making the motive of $G_{\bf e}$ explicit. Defining polynomials $n_{\bf d}$ by the same recursion with $\mathbb{L}$ replaced by $x$, and starting with $n_{0}=1$, we obviously have $n_{\bf d}\in\mathbb{Z}[x^{\pm 1}]$, and $n_{\bf d}(\mathbb{L})=[N_{\bf d}]$. By the definition of the Euler form, the $\mathbb{L}$-exponent in the recursion is always nonnegative, thus we have $n_{\bf d}(x)\in\mathbb{Z}[x]$.∎
@@
which equals 00 for ${\bf d}\not=0$ by Theorem 2.2.∎
@@
=$\sum_{s\geq 0}(-1)^{s}\cdot\sum_{{\bf d}^{1},\ldots,{\bf d}^{s}\not=0}\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\frac{t^{\bf d}}{(\mathbb{L})_{{\bf d}^{1}}\cdot\ldots\cdot(\mathbb{L})_{{\bf d}^{s}}}.$
Comparing coefficients of $t^{\bf d}$ on both sides and rewriting $[G_{\bf d}]$ using the Pochhammer symbol, we arrive at the claimed formula.∎
@@
fulfills $x^{\bf d}\cdot x^{\bf e}=x^{{\bf d}+{\bf e}}$; consequently, we do not have to work with twisted multiplications of formal series in the symmetric case. We extend the base ring of motives by adjoinging a square root of $\mathbb{L}$. This allows us to define motivic Donaldson–Thomas invariants (see [14] for an exposition) by the identity
@@
Polynomiality of the Donaldson-Thomas invariants allows us to formally interchange $\mathbb{L}$ and $\mathbb{L}^{-1}$, and thus we have
@@
Combining these identities, the claim follows.∎
@@
$\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i,j}\sum_{k}r_{ij}\delta_{k\in K_{i}}\delta_{k\in K_{j}})=$
$=\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i}\sum_{k}r_{ii})=\sum_{i}r_{ii}{d_{i}\choose 2}+\sum_{i<j}r_{ij}d_{i}d_{j}.$
All of the corresponding summands have the same sign $(-1)^{\dim{\bf d}}$, thus there are no cancellations in the summation of these top degree terms. The number of these summands equals the number of decompositions of $\{1,\ldots,\dim{\bf d}\}$ into parts $K_{i}$ for $i\in Q_{0}$, which is the quotient of factorials of the claimed formula. Adding the contribution $\sum_{i}{d_{i}\choose 2}$ of the prefactor in the above formula for $n_{\bf d}(x)$ to the degree, we arrive at the claimed formula for the leading term.∎
Remark: A similar proof appears in [13, Theorem 4.1]. Refining the argument there shows that, in case $Q$ is symmetric, the variety $N_{\bf d}$ is equidimensional of dimension
@@
    $t_{i_{n}}^{d_{i_{n}}}\cdot\ldots\cdot t_{i_{1}}^{d_{i_{1}}}=\mathbb{L}^{-\sum_{k}d_{i_{k}}(d_{i_{k}}-1)/2}t^{\bf d}.$
@@
    $\sum_{\bf d}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}}=\prod_{k=1}^{n}\sum_{d\geq 0}\frac{\mathbb{L}^{d(d-1)/2}t_{i_{n+1-k}}^{d}}{(\mathbb{L})_{d}}$
@@
    $(\sum_{\bf d}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}})^{-1}=\prod_{k=1}^{n}\sum_{d\geq 0}\frac{(-t_{i_{k}})^{d}}{(\mathbb{L})_{d}}=$
@@
    It is left as an exercise to the reader that, for general $d$, a monomial $\mathbb{L}^{a}$ appears in $[N_{d}]$ with non-zero coefficient if and only if $a=km+l$ for $d-1\leq k\leq{d\choose 2}$ and (${d\choose 2}-k\leq l\leq{d\choose 2}$.
@@ 4.1. Introduction
$I_{v}=[\sum_{i=1}^{v-1}d_{i}+1\ldots\sum_{i=1}^{v}d_{i}]~{}.$
Note that we need to fix an enumeration of the vertices of $Q$, but the following constructions will be essentially independent of this choice.
@@
If $R$ is unstable, there is a unique coweight $s(R)\in\mathbb{Q}^{d}$ (see [9, Lemma 6.2]) with this property and such that the norm $|s(R)|=s_{1}^{2}+\dots+s_{d}^{2}$ is minimal. We will call such a coweight minimal.
@@
d) We denote by $S_{\mathbf{d}}=S_{d_{1}}\times\dots\times S_{d_{n}}$ the Weyl group of $G_{\mathbf{d}}$.
@@
(see [8, Section 2.2]) If $R$ is a saturated subset of $\Pi$, we have a corresponding saturated subspace $X_{R}$ of $N_{\mathbf{d}}(Q)$ by taking
@@
By [5] there is a bijection between the $G_{\mathbf{d}}$-conjugacy classes of saturated subspaces of $N_{\mathbf{d}}(Q)$ and the conjugacy classes of saturated subsets $R\subset\Pi$ under the action of the Weyl group $S_{\mathbf{d}}$. This correspondence assigns to $R$ the subspace $X_{R}$, and to a saturated subspace the set of nonzero weights of its elements. Hence, the number of Hesselink strata of $N_{\mathbf{d}}(Q)$ is at most the number of conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group.
We construct a bijection between the conjugacy classes of saturated subsets $R\subset\Pi$ under the Weyl group and coweights satisfying certain additional properties. Consider a coweight $s=(s_{1},\dots,s_{d})\in\mathbb{Q}^{d}$ for $G_{\mathbf{d}}$. Then there is a unique partition of $[1\ldots d]$ into a minimal number of pairwise disjoint subsets $\tilde{J}$, such that the $s_{j}$ for $j\in\tilde{J}$ form an ordered sequence increasing by 00 or $1$ in each step after sorting. We call such a sequence a segment and denote it by $J$. After introducing this notation we can now state definitions for the required conditions on coweights.
@@
Since the property of being minimal can be hard to verify, our goal is to extend the index set in such a way, that the new summands in the equation (1) are equal to zero, and it is easy to verify if a string belongs to the new index set. The property of being balanced will play an essential role and the following lemma portrays the first step [8], see also [5].
@@
A minimal coweight $\mu=(\mu_{1},\dots,\mu_{d})\in\mathbb{Q}^{d}$ is balanced.
@@
Suppose $\mu=(\mu_{1},\dots,\mu_{d})$ is not balanced. Then there exists a segment $J$ such that $\sum_{i\in\tilde{J}}\mu_{i}\neq 0$. By definition of the segments, for every $\mu_{j}\in J$ and $\mu_{k}\notin J$ we have $|\mu_{j}-\mu_{k}|<1$ or $|\mu_{j}-\mu_{k}|>1$. Consequently we can choose an $\epsilon>0$ such that these inequalities also hold for $\mu_{j}+\epsilon$ and $\mu_{j}-\epsilon$. By shifting all entries of $J$ by $\epsilon$ or $-\epsilon$, the distance between two entries in the segment does not change. In this way we construct another string $\mu^{\prime}$ such that $R(\mu)=R(\mu^{\prime})$. Without loss of generality let $\sum_{i\in\tilde{J}}\mu_{i}=\alpha>0$ and we define $\mu^{\prime}$ by
$\mu_{i}^{\prime}=\begin{cases}\mu_{i}&\text{if }i\notin\tilde{J}\\ \mu_{i}-\epsilon&\text{if }i\in\tilde{J}~{}.\end{cases}$
We denote the cardinality of $\tilde{J}$ by $n$ and obtain:
$|\mu^{\prime}|-|\mu|$ $=\sum_{i=1}^{d}(\mu^{\prime}_{i})^{2}-\sum_{i=1}^{d}(\mu_{i})^{2}=\sum_{\tilde{J}}(\mu_{i}-\epsilon)^{2}-\sum_{\tilde{J}}(\mu_{i})^{2}$
$=\sum_{\tilde{J}}-2\mu_{i}\epsilon+\epsilon^{2}=-2\alpha\epsilon+n\epsilon^{2}~{}.$
@@
By choosing $\epsilon<\frac{2\alpha}{n}$, we obtain $|\mu^{\prime}|<|\mu|$. Accordingly, $\mu$ is not minimal. The case $\sum_{i\in\tilde{J}}\mu_{i}<0$ follows analogously with
$\mu_{i}^{\prime}=\begin{cases}\mu_{i}&\text{if }i\notin\tilde{J}\\ \mu_{i}+\epsilon&\text{if }i\in\tilde{J}~{}.\end{cases}$
@@
We can thus extend the index set of the Hesselink strata using balanced and dominant coweights. The problem of detecting which strata are non-empty remains. To this effect, for every coweight $s$, we define a tuple $(Q_{s},{\bf d}_{s},\theta_{s})$, consisting of a quiver, a dimension vector and a stability, as follows:
@@
$\{(v,j)\;|\;v\in Q_{0}\text{ and }0\leq j\leq k_{i}\}~{},$
where $p_{i}+k_{i}$ is the last entry in $J_{i}$. In $Q_{i}$ the number of arrows from $(v,k)$ to $(v^{\prime},k+1)$ equals the number of arrows from $v$ to $v^{\prime}$ in $Q$. The dimension vector ${\bf d}_{i}$ of $Q_{i}$ is given by $(a_{i,k}^{v})_{v,k}$.
@@
A stability on $Q_{s}$ is defined by $\theta_{s}=(m_{v,j}^{(i)})_{i,v,j}$, where $m_{v,j}^{(i)}=-z\cdot(p_{i}+j)$ and $z$ is the least common multiple of all denominators of the rational numbers $p_{i}$ with $1\leq i\leq u$.
@@
The following theorem of [8] summarizes the main properties relating the representation spaces of level quivers and the Hesselink stratification:
@@
Consider $|Q_{1}|$-tuples of $d\times d$-matrices. Then $V_{s}$ is given as the affine space of such tuples, where the matrix corresponding to an arrow $\alpha:v\rightarrow v^{\prime}$ has arbitrary entries at a position $(j,i)$ if $i\in I_{v}$, $j\in I_{v^{\prime}}$ and $s_{j}-s_{i}\geq 1$, and zero entries everywhere else (compare with Definition 4.1). The group $P_{s}$ is given by a single invertible $d\times d$-matrix with zero entries everywhere except possibly at a position $(j,i)$ if $i,j\in I_{v}$ for some vertex $v\in Q_{0}$ and $s_{j}-s_{i}\geq 0$. For the affine subspace $B_{s}\subset V_{s}$ obtained by additionally requiring zero entries at all positions $(j,i)$ except the ones where $s_{j}-s_{i}=1$, we have